The state parks department is considering the purchase of of a marsh to form a bird sanctuary. Bird watching is valued at $2 per hour, the state interest rate is 5% and perpetual life is assumed. Viewing hours is estimated at 400,000 per year. what is the EAW for the sanctuary? what fraction of the EAW is due to the due to the expected increases?
Answer #1

Value of bird watching per hour = $2

No. of hours of bird watching per year = 400000

So, annual revenue from bird watching = 2*400000 = $800000

R = 5%

So,

Total worth = 800000/5% = 16000000 ( since it is a case of perpetuity)

EAW = 16000000*5%

EAW = $800000